The time to do a certain job varies inversely as the
number N of people working. It takes 5 hr for 24 people
to wash and wax the floors in a building. How long
would it take 36 people to do the job?

Answer by stanbon(75887)

The time to do a certain job varies inversely as the
number N of people working.
t = k/N
--------------------
Solve for "k" using::"It takes 5 hr for 24 people
to wash and wax the floors in a building".
5 = k/24
k = 120
---------
How long would it take 36 people to do the job?
Equation:
t = 120/N
---
t = 120/36
time = 3 1/3 hrs
-----------------------
Cheers,
Stan H.
--------------------